OpenStudy (anonymous):

how do we subtract 8 1/3 and 7/9 ?

OpenStudy (anonymous):

First you need to convert mixed fraction 8 1/3 into an improper fraction. So, 8 + 1/3 = ?

OpenStudy (anonymous):

like this (8 1/3) -(7/9)

OpenStudy (anonymous):

further gt bro will explain to u

OpenStudy (anonymous):

I don't get it yet

OpenStudy (anonymous):

Adding 8 1/3 and 7/9 as is won't work. You need both numbers in the same form. 8 1/3 is a mixed fraction. 7/9 is a simple fraction. So, you need to convert the first one to the same form. 8 1/3 is nothing but 8 + 1/3. Write it in a fraction form. If you can't do that, I suggest you go back and study that part first.

OpenStudy (amistre64):

8 1/3 - 7/9 8 + 1/3 - 7/9 = N 3(8 + 1/3 - 7/9 = N) 24 + 1 -7/3 = 3N 3(24 + 1 -7/3 = 3N) 75 -7 = 9N 68 = 9N thats how id do it

OpenStudy (anonymous):

What does the N stand for?

OpenStudy (amistre64):

the N is just a placeholder for some "N"umber. Its there so i can algebra it into submission

OpenStudy (amistre64):

in other words, solve for the Number :)

OpenStudy (anonymous):

things r very simple but we dont try to apply logic so it looks difficult

OpenStudy (anonymous):

amistre64 - this is a simple arithmetic problem. no need to confuse the person with algebraic stuff. :-)